Meeting Information

  • 1930 UT (1330 MDT) -- NOTE: Changing!
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News

  • New Meeting time: Tuesdays at 1000 MDT in AOC 317; on weeks with a Monday holiday, e.g., July 6, Sept 7, Dec 28(?) - it will start at 1100 after the coordination meeting.
    • No meeting next week due to the summer school.
  • Need to enhance the commissioning testing effort; still relying on small groups to drive commissioning efforts. Refocus meeting to review planning, status and results.
  • Staff Travel:
    • Emmanuel: 21 May - 7 June
    • Gustaaf: 1 - 4 June; 18 June; 21 June
    • Eric: 4 - 19 June
  • Key Dates:
    • Synthesis Imaging Summer School: 8-15 June
    • PASEO Meeting: 15-16 July
    • VLA 30th anniversary: 12 August

Critical Commissioning review/reports

System Verification/Hardware Commissioning:

Testing and characterization of the array based on the design goals.

  • News from the weekend+: Many issues. Lost most of Fri-Sat due to a Lustre disk issue; the problem unfortunately masked itself as a host of other problems which were meticulously worked through until the real culprit was exposed. Monday testing resulted in complete reboot of StB/BB/etc. Issues with missing scans/missing visibilities tied to correlator configuration changes. Last two days have fallen back to low frequency single config files while the issues were being diagnosed/tested. Several test programs run - need to discuss results (below).

  • Antenna: Perley
    • EA10 passed operations acceptance testing (last of the VLA -- Matt Gardiner)
  • Receiver: Perley
    • Rick: Pol'n tests (receiver rotation) to be scheduled
  • Array: Perley
    • Switched power tests
    • Doppler tracking updates
  • Correlator: Rupen
    • Missing Scans (Joe et al)
    • Lost Fringes (Vivek et al)
    • In-band frequency change results (Miriam) Report
    • Wide Band testing results: Dhawan, Greisen, Krauss, Owen, Perley, Brogan, Hunter, Chomiuk, Marvil
    • Visitors
      • Sonja Vrcic: 21 June - ?
      • Brent Carlson: x July - ?

Observing Band Checkout

Encompasses principally core operation at those frequencies followed by detailed calibration strategies, flagging.

Summary testing status by band: https://staff.nrao.edu/wiki/bin/view/EVLA/EVLAObservingBandCheckouts

  • Low Frequency (L/S/C): Momjian
    • Status: Operating; users should include one of the canonical flux calibrators in SB.
    • Upcoming:
      • Need switched power
        • Develop set of observations to check Tcal/Tsys/SEFD across bands
    • Links: ECSV-38
  • Mid Frequency (X/Ku): Dhawan
    • Status: Operating; users should include one of the canonical flux calibrators in SB.
    • Upcoming:
      • Need switched power
        • Develop set of observations to check Tcal/Tsys/SEFD across bands
    • Links
  • High Frequency (K/Ka/Q): Sjouwerman/Claussen
    • Status:
      • Operating; users should include one of the canonical flux calibrators in SB
      • Reference pointing in C or X
      • No observing at Ka < 32 GHz
    • Upcoming:
      • Need switched power
        • Develop set of observations to check Tcal/Tsys/SEFD across bands
    • Links: ECSV-30

Observing Mode Support:

Testing and development within broad observing modes distinguished by their science goals (continuum, spectral line, polarization).

  • Spectral Line Modes (Ott/Goss)
  • Polarization Modes (Myers)
    • Testing
      • C (ECDATA-12): Ready
      • L (ECDATA-42): Ready
      • Combined Low-High Pol checkout test (TPOL0003_sb1378180): Report
      • Post-processing (AIPS): AH1004 data checkout
    • Links: ECSV-29

  • Advanced: Phased Array (Lazio/Deller)
  • Advanced Planetary:
    • Status:
      • Executor: Objects not supported in executor.
      • SDM: Ephemerides/motion not supported.
  • Advanced: Solar
    • Status: Planning/prototype testing ongoing; working with L, Ku (special Rx setups in these two bands). TP tests coming.
  • Advanced: Pulsar (Brisken)
  • Advanced: VLBI (Walker/Brisken)
  • Advanced: Astrometry (Mioduszewski)

Observing Support:

Testing and development of the end-to-end capability of the array to support observations to the community (Documentation, OPT, Scheduling, Instrument models, Calibrator survey/models DB, Rapid response observing, demo science, post-processing).

  • Meta summary
    • Multi-config SBs in regular use; note 'overhead' in configuration changes; Must add 30 seconds to each configuration change - Gustaaf's group will check projects for this until relaxed, in particular, reference pointing must expand to 3 minutes..

Mode OPT Scheduling Executor/Corr Calibration strategy Archive CASA AIPS Data Distribution Other Overall Status
  wide band modes available only in webtest OST (reconciling status database)

OPT selected limits not possible
Running 20 Hz
18-May CM version

21-May CBE version

    - needs check on short pol'n runs Need decision on large (i.e., RSRO) data set distribution media   Operational for OSRO1/OSRO2
30s overhead with OSRO configuration changes
RSRO beginning

  • Software Systems
    • OPT
    • Scheduling
    • Executor
    • Corr (CBE/CM)
    • Archive
    • Post-processing

  • RSRO
    • Status: RSRO Readiness (Cycle1) Master List
    • Resident Schedule: Posted on home page of twiki now!
      • 01 Jan - 16 May: Adam Deller (AC982: RSRO-4): 300A
      • 31 Jan - 15 Jul: Laura Chomiuk (AS1015/AK726: RSRO-11/RSRO-6): 200C
      • 04 Feb - 01 Jun: Jacqueline van Gorkom (AV318: RSRO-13): 200D
      • 25 Feb - 01 Jun: James Miller-Jones (AM1014: RSRO-9): 200E
      • 02 Mar - 07 Jun: Manuel Aravena (AA330: RSRO-1): 300G
      • 16 Mar - 08 Apr: Dave Green (AB1345: RSRO-2): 300G
      • 20 Apr - 30 Jun: Crystal Brogan (AB1346: RSRO-3): 200B
      • 20 Apr - 30 Jun: Todd Hunter (AB1346: RSRO-3): 200B
      • 30 Apr - 15 Jul: Joseph Lazio (AC982: RSRO-4): 266
      • 07 Jun - 16 Sep: Laura Perez (AC982: RSRO-4): 200E
      • 28 Jun - 05 Oct: Volker Heesen (AH1006: RSRO-5): 200G
      • 01 Jul - 31 Oct: Russ Taylor (AT374: RSRO-12): 300G
      • 15 Sep - 15 Dec: Adam Leroy (AL746: RSRO-7): 200C
      • 01 Oct - 01 Jan: Hendrik Linz (AC982: RSRO-4): 200C
      • 01 Jan - 01 Apr: Jane Greaves (AC982: RSRO-4): 300G
      • 02 Jan - 02 Apr: Stuart Corder (AC982: RSRO-4): 300G
    • Current Residents
      • Manuel Aravena (POC: Chris Carilli); area(s): High frequency checkout (reference pointing tests); OSRO data validation
        • This week:
      • Crystal Brogan (POC: Claire Chandler); area(s): High frequency observing; demo science
        • This week:
      • Laura Chomiuk (POC: Miller Goss; area(s): Correlator checkout; Spectral line commissioning; OSRO data validation
        • This week:
      • Adam Deller (POC: Michael Rupen); area(s): phased array system design and implementation
        • This week:
      • Todd Hunter (POC: Claire Chandler; area(S): science support software interfaces for users and high frequency observing; band/polarization checkout.
        • This week:
      • Joe Lazio (POC: Claire Chandler): Low frequency observing.
        • This week:
    • ---
    • Dave Green (POC: Sanjay Bhatnagar): area(s): algorithm tests for wideband imaging (currently away).
    • James Miller-Jones (POC: Claire Chandler; area(s): demo science use cases
    • Jacqueline van Gorkom (POC: Miller Goss; area(s): Correlator checkout; spectral line commissioning)

  • Documentation

Advanced Science Support:

These are the testing/development areas needed to fulfill the full capabilities of the EVLA along with additional observing modes for other branches of the community.

  • Astrometry (Mioduszewski)
    • Status: Wait for A array(?)

AOB

Actions

  • [Rupen]: Develop daily/post-maintenance day checklists for the system. To be completed as part of stress tests.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley]: Develop set of frequency accuracy tests (X band is complete).
  • [Perley]: Perform a frequency sweep test at different bands to provide to engineers for review (Q2).
  • [Wrobel/Butler]: Make use of API criteria consistent throughout tools (OPT->OST). List of actions organized. In process.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Subband 0 problems: Being reviewed on Wednesday morning at the WIDAR meeting. Butler to report.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: Occasional zero-amplitude visibilities (needs software fix - what level). UVLOD and FITLD can handle this in AIPS; flagdata clip is effective in CASA; still need to understand and fix at the source.
      • Low Priority [TBD]: Occasional pure-noise scans.
      • Low Priority [Rupen/Pokorny]: Delay clunking; software fix pending.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.

-- JosephMcMullin - 01 Jun 2010
Topic revision: r3 - 2010-06-02, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback